# The Topology of Rayleigh-Levy Flights in Two Dimensions

### v1

@@
11institutetext: Asia Pacific Center for Theoretical Physics, Pohang 37673, Republic of Korea 22institutetext: Department of Physics, POSTECH, Pohang 37673, Republic of Korea 33institutetext: Université Paris-Saclay, CNRS, CEA, Institut de physique théorique, 91191, Gif-sur-Yvette, France 44institutetext: Institut d’Astrophysique de Paris, 98 bis Boulevard Arago, F-75014 Paris, France 55institutetext: Kyung Hee University, Dept. of Astronomy & Space Science, Yongin-shi, Gyeonggi-do 17104, Republic of Korea
@@
Reginald Christian Bernardo , 11    Stephen Appleby , 1122    Francis Bernardeau , 3344    Christophe Pichon , Corresponding author: pichon@iap.fr334455
@@ 1 Introduction
Previous work Bernardeau (2022); Bernardeau & Pichon (2024) derived from first principle one- and two-point statistics for the critical points of smoothed sampled realisation of these Rayleigh-Lévy flights in one to three dimensions. They provided closed-form expressions for Euler counts and their clustering correlations in the mean-field limit, describing their long-separation behaviour. They also presented quadratures for calculating critical point number densities. Their result was validated by Monte Carlo simulations in one dimension only. Following Bernardeau & Pichon (2024), the present study focuses on validating in two dimensions both the density field probability distribution and associated Minkowski functionals. We restrict ourselves to expansion beyond the mean field for the density PDF and iso-field perimeter, $W_{1}$, and leave the exploration of quantities such as Euler characteristics or extrema counts in 2D to future work.
@@ 2.1 Basics
The Rayleigh-Lévy flight cumulative distribution function is given by
[…]
@@
where $n$ is the number density of points and $\psi(k)$ is the Fourier transform of $f_{1}({\mathbf{r}})$, the density of the subsequent point at position ${\mathbf{r}}$ (first descendant). For a Rayleigh-Lévy flight, this is given by
$f_{1}({\mathbf{r}})=\frac{\alpha\Gamma(D/2+1)}{2\pi^{D/2}}\frac{\ell_{0}^{\alpha}}{|{\mathbf{r}}-{\mathbf{r}}_{0}|^{D+\alpha}}\,.$ (3)
The large distance correlation ($kl_{0}\ll 1$ or $1-\psi(k)\sim(\ell_{0}k)^{\alpha}$) behaves like
@@
This is subject to constraints ${\alpha<D}$ and ${0<\alpha<2}$; the first is required to avoid a logarithmic divergence at ${k=0}$ and the second comes from the expected behavior of $\psi(k)$ at low $k$, ${\psi(k)=1-(l_{0}k)^{\alpha}+{\cal O}(k^{2})}$.
@@
$\varphi(\lambda_{1},\dots,\lambda_{n})={\hskip-5.69046pt}\sum_{p_{1},\dots,p_{n}}\langle\rho_{1}^{p_{1}}\dots\rho_{n}^{p_{n}}\rangle_{c}\,\frac{\lambda_{1}^{p_{1}}}{p_{1}!}\,.$ (5)
@@
A recipe for the CGF is the remaining piece. This is provided by the hierarchical tree model (Appendix B of Bernardeau & Pichon (2024)). In this case, the CGF can be approximated via the set of equations,
@@
where $\varphi(\lambda_{1},\dots,\lambda_{n})$ is the CGF, $\tau({\mathbf{x}})$ is the rooted-CGF (r-CGF), ${\cal C}_{j}({\mathbf{x}})$ is a window function in the $j$th cell, and $\xi({\mathbf{x}},{\mathbf{x}}^{\prime})$ is the two-point correlation function. This will be used in the next section to study the properties of Rayleigh-Lévy flights in the so-called mean field limit.
@@ 2.2 Mean field
[…]
@@
[…]
@@
For a Gaussian window function,
${\cal C}_{0}(r)\equiv{\cal C}_{0}(x){\cal C}_{0}(y)=e^{-r^{2}/2}/(2\pi)\,,$ (15)
similarly to the 1D case, the eigenfunctions of the hamiltonian of a 2D harmonic oscillator can be utilized. In general those solutions will be polynomials in both $x$ and $y$ multiplied by ${\cal C}_{0}({\mathbf{x}})$. The overall degree of the polynomial gives the total order [^1] The total order $n_{t}$ corresponds to the energy level $(1+n_{t})E_{0}$ of the associated harmonic oscillator. $n_{t}$. Alternatively one can use angular coordinates, $(r,\theta)$, where $x=r\cos(\theta),\ y=r\sin(\theta)$, to describe the solution of a 2D harmonic oscillator. This leads to a choice of basis of the form
@@
Figure 3: Left: Two-dimensional Lévy flight probability distribution function and right: difference with respect to ‘mean field’ (MF) with parameters $\alpha=3/2$, $\xi_{0}=0.08,0.88,2.47$ (top-bottom panels), $L=4000\times\Delta$ (box width), $N_{\rm pts}=40\times 2000^{2}$ (number of steps), $\Delta=0.25$ (resolution), and $R=1$ (smoothing scale). The error bars show the standard deviation in the left panels and the error on the mean for $100$ realisations in the right panels. The points the sample means. The corresponding standard deviations of the fields are given by $\sigma_{0}=\sqrt{\xi_{0}}=0.28,0.94,1.57$ (top-bottom). The BMF prediction agrees well with the simulations even in the non-perturbative, non-Gaussian regime.
@@
$\varphi_{4}(\lambda)$ $=-\frac{\lambda\left(225\lambda^{2}\xi_{0}^{2}-59680\lambda\xi_{0}+1048576\right)}{80\lambda^{3}\xi_{0}^{3}-25569\lambda^{2}\xi_{0}^{2}+583968\lambda\xi_{0}-1048576}\,.$
Appendix A.2 spells out higher CGFs, up to $\varphi_{6}(\lambda)$, together with the skewness ${\cal S}_{3}$ derived within the BMF formalism up to 3rd order (terms in Eq. 18) compared with the exact value.
The CGFs in the BMF formalism can also be obtained for an arbitrary power law correlation function [$\xi(r)=1/r^{\beta}$]. The corresponding CGFs can be written analytically as
@@
In the next section, we compare the mean field and beyond mean field predictions with simulations through the lens of one-point statistics; the PDF of the density, the Minkowski functional $W_{1}$ and the Euler characteristic. The latter two are defined as follows
@@
×$\times\frac{\partial}{\partial\lambda_{\rho}}\left[\frac{\partial\varphi(\lambda_{\rho},\lambda_{g})}{\partial\lambda_{g}}\exp\left(\varphi(\lambda_{\rho},\lambda_{g})\right)\right]\,.$ (30)
@@ 4 Results
The Lévy flights are generated starting from an initial point in the center of a box of area $L\times L$. The flight is moved a distance $\ell$ from the previous point, in a direction drawn from a uniform angle $0\leq\theta<2\pi$. The $i^{\rm th}$ point in the flight is given by $\bf{x}_{i}={\bf x}_{i-1}+(\ell_{i}\cos\theta_{i},\ell_{i}\sin\theta_{i})$. Periodic boundary conditions are enforced, and ${\bf x}_{i}$ is adjusted modulo $L$ if the flight length is greater than $L$. Each point in the flight is binned into a regular lattice using a nearest-neighbour scheme, and we use resolution $\Delta=0.25$ and number of pixels per side in the box $N_{\rm pix}=4000$, so that the box size is $L=N_{\rm pix}\times\Delta$. We fix the number of Lévy flight steps as $N_{\rm pts}=40\times 2000^{2}$, $\alpha=3/2$ and vary $\ell_{0}$ to obtain fields with different $\xi_{0}$ values. Once the Lévy flight is binned into a regular lattice, we fast Fourier transform the field, apply a Gaussian smoothing kernel $W(kR)=e^{-k^{2}R^{2}/2}$, and inverse transform. We fix $R=1$ to match the window function used in the previous section.
@@
Before presenting the results, we first comment on a numerical systematic which hinders our ability to exactly reproduce the theoretical expectation values constructed in the previous section. In Figure 2 (top panel) we present the two-point correlation function of a set of $N_{\rm real}=100$ numerically generated Lévy flights with $\alpha=3/2$ and $\ell_{0}=0.02$ (red points/error bars). The expected large separation behaviour of the flight is $\zeta(r)\sim r^{\alpha-D}$ with $D=2$; this expected behaviour is presented as a blue dashed line. However, we find that the simulated Lévy flights present slightly steeper power law behaviour (solid blue line), corresponding to an effective $\alpha=1.43$. This is a finite box effect; each step in the flight must be smaller than $\sim L$ due to the imposition of periodic boundary conditions. This leads to an excess probability of finding flight points at smaller separations, and leads to a systematic difference between theoretical expectation values and our numerical results [^2] Or alternatively one could model finite volume effects by replacing in equation (2) the integral over $\mathbf{k}$ by a sum over discrete values.. The problem will be exacerbated for larger values of $\alpha$, as the correlation function should become flat as $\alpha\to D$, which is a scenario that cannot be realised for a finite box. In what follows, for all summary statistics we present the expected theoretical results based on $\alpha=3/2$ and the effective result assuming $\alpha=1.43$. For mean field predictions, there is no $\alpha$ dependence.
@@ 5 Conclusion
Our results highlight the versatility of Rayleigh-Lévy flights, offering insights not only for astrophysical contexts but also for broader interdisciplinary applications.

### v2

@@
11institutetext: Max Planck Institute for Gravitational Physics (Albert Einstein Institute), Hannover 30167, Germany 22institutetext: Asia Pacific Center for Theoretical Physics, Pohang 37673, Republic of Korea 33institutetext: Department of Physics, POSTECH, Pohang 37673, Republic of Korea 44institutetext: Université Paris-Saclay, CNRS, CEA, Institut de physique théorique, 91191, Gif-sur-Yvette, France 55institutetext: Institut d’Astrophysique de Paris, 98 bis Boulevard Arago, F-75014 Paris, France 66institutetext: Kyung Hee University, Dept. of Astronomy & Space Science, Yongin-shi, Gyeonggi-do 17104, Republic of Korea
@@
Reginald Christian Bernardo ,    Stephen Appleby ,    Francis Bernardeau ,    Christophe Pichon , Corresponding author: pichon@iap.fr
@@ 1 Introduction
Previous work Bernardeau (2022); Bernardeau & Pichon (2024) derived from first principle one- and two-point statistics for the critical points of smoothed sampled realisation of these Rayleigh-Lévy flights in one to three dimensions. They provided closed-form expressions for Euler counts and their clustering correlations in the mean-field limit, describing their long-separation behaviour. They also presented quadratures for calculating critical point number densities. Their result was validated by Monte Carlo simulations in one dimension only. Following Bernardeau & Pichon (2024), the present study focuses on validating in two dimensions using the Minkowski Functionals – a set of summary statistics comprising the probability density function of the density field, the total perimeter length of iso-field contours, and the Euler characteristic, which together characterize the non-Gaussianity of the field’s one-point distribution. We restrict ourselves to expansion beyond the mean field for the density PDF and iso-field perimeter, $W_{1}$, and leave the exploration of quantities such as Euler characteristics or extrema counts in 2D to future work.
@@ 2.1 Basics
A Rayleigh-Lévy flight is a point distribution for which each successive point is a step from the previous, in a random direction and with a length drawn from a cumulative probability distribution function of the form
${\rm CDF}(l)=\cases{0}\,&,\ \ \ \ l<l_{0}\\ 1-\left(l_{0}/l\right)^{\alpha}\,&,\ \ \ \ l\geq l_{0}\,.{}$ (1)
@@
where $n$ is the number density of points and $\psi(k)$ is the Fourier transform of $f_{1}({\mathbf{r}})$, the density of the subsequent point at position $\mathbf{r}$} (first descendant of a given point at position $\mathbf{r}_{0}$}). For a Rayleigh-Lévy flight, this is given by
$f_{1}({\mathbf{r}})=\frac{\alpha\Gamma(D/2+1)}{2\pi^{D/2}}\frac{\ell_{0}^{\alpha}}{|{\mathbf{r}}-{\mathbf{r}}_{0}|^{D+\alpha}}\,$ (3)
for $|{\mathbf{r}}-{\mathbf{r}}_{0}|>l_{0}$. The large distance correlation ($kl_{0}\ll 1$ or $1-\psi(k)\sim(\ell_{0}k)^{\alpha}$) behaves like
@@
This is subject to constraints $\alpha<D$} and $0<\alpha<2$}; the first is required to avoid a logarithmic divergence at k=0{k=0} and the second comes from the expected behavior of $\psi(k)$ at low $k$, $\psi(k)=1-(l_{0}k)^{\alpha}+{\cal O}(k^{2})$}. [^1] This is expected assuming that the correlation function will not exhibit divergent behavior at large distances. The asymptotic behavior can be obtained starting with an ansatz $1-\psi(k)=(l_{0}k)^{\alpha}$ and performing (2) bearing in mind that the correlation function must not diverge at $k=0$ and large distances $kl_{0}\ll 1$.
@@
$\varphi(\lambda_{1},\dots,\lambda_{n})={\thinspace}\sum_{p_{1},\dots,p_{n}}\langle\rho_{1}^{p_{1}}\dots\rho_{n}^{p_{n}}\rangle_{c}\,\frac{\lambda_{1}^{p_{1}}}{p_{1}!}\,.$ (5)
@@
The remaining piece required to estimate the PDF of the density field is a recipe for the the CGF. This is provided by the hierarchical tree model (Appendix B of Bernardeau & Pichon (2024)). In this case, the CGF can be approximated via the set of equations,
@@
where $\varphi(\lambda_{1},\dots,\lambda_{n})$ is the CGF, $\tau({\mathbf{x}})$ is the rooted-CGF (r-CGF), ${\cal C}_{j}({\mathbf{x}})$ is a window function in the $j$th cell, and $\xi({\mathbf{x}},{\mathbf{x}}^{\prime})$ is the two-point correlation function [^2] We define a ‘cell’ as an area occupied by the smoothing window function, which has typical size $\sim R_{G}$ where $R_{G}$ is the associated smoothing length. We fix $R_{G}=1$ throughout this work.. This will be used in the next section to study the properties of Rayleigh-Lévy flights in the so-called mean field limit.
@@ 2.2 Mean field
$\tau_{i}\!=\!\sum_{j}\lambda_{j}\,\left(1\!+\!{\tau_{j}}/{2}\right)\!\int\!{d{\mathbf{x}}}\!\int\!\!{d{\mathbf{x}}^{\prime}}\,{\cal C}_{i}({\mathbf{x}})\,{\cal C}_{j}({\mathbf{x}}^{\prime})\,\xi({\mathbf{x}},{\mathbf{x}}^{\prime})\,.$ (9)
@@
$\xi_{0}=\int{d{\mathbf{x}}}\,{d{\mathbf{x}}^{\prime}}\,{\cal C}_{1}({\mathbf{x}})\,{\cal C}_{1}({\mathbf{x}}^{\prime})\,\xi({\mathbf{x}},{\mathbf{x}}^{\prime})\,\,.$ (11)
@@
For a Gaussian window function, we have
${\cal C}_{0}({\bf r})\equiv{\cal C}_{0}(x){\cal C}_{0}(y)=e^{-r^{2}/2}/(2\pi)\,$, (15)
where $r^{2}=|{\bf r}|^{2}$. Similarly to the 1D case, the eigenfunctions of the Hamiltonian of a 2D harmonic oscillator can be utilized. In general those solutions will be polynomials in both $x$ and $y$ multiplied by ${\cal C}_{0}({\mathbf{x}})$. The overall degree of the polynomial gives the total order [^3] The total order $n_{t}$ corresponds to the energy level $(1+n_{t})E_{0}$ of the associated harmonic oscillator. $n_{t}$. Alternatively one can use angular coordinates, $(r,\theta)$, where $x=r\cos(\theta),\ y=r\sin(\theta)$, to describe the solution of a 2D harmonic oscillator. This leads to a choice of basis of the form
@@
Figure 3: Left: Two-dimensional Lévy flight probability distribution function and right: difference with respect to ‘mean field’ (MF) with parameters $\alpha=3/2$, $\xi_{0}=0.08,0.88,2.47$ (top-bottom panels), $L=4000\times\Delta$ (box width), $N_{\rm pts}=40\times 2000^{2}$ (number of steps), $\Delta=0.25$ (resolution), and $R=1$ (smoothing scale). The error bars show the standard deviation in the left panels and the error on the mean for 100100 realisations in the right panels. The points the sample means. The corresponding standard deviations of the fields are given by $\sigma_{0}=\sqrt{\xi_{0}}=0.28,0.94,1.57$ (top-bottom). The BMF prediction agrees well with the simulations even in the non-perturbative, non-Gaussian regime.
@@
$\varphi_{4}(\lambda)$ $=-\frac{\lambda\left(225\lambda^{2}\xi_{0}^{2}-59680\lambda\xi_{0}+1048576\right)}{80\lambda^{3}\xi_{0}^{3}-25569\lambda^{2}\xi_{0}^{2}+583968\lambda\xi_{0}-1048576}\,$,
where the index $i$ on $\varphi_{i}$ indicates the BMF order. Appendix A.2 spells out higher CGFs, up to $\varphi_{6}(\lambda)$, together with the skewness ${\cal S}_{3}$ derived within the BMF formalism up to 3rd order (terms in Eq. 18) compared with the exact value.
The CGFs in the BMF formalism can also be obtained for an arbitrary power law correlation function $\xi(r)=1/r^{\beta}$]. The corresponding CGFs can be written analytically as
@@
In the next section, we compare the mean field and beyond mean field predictions with simulations through the lens of one-point statistics; specifically the Minkowski functionals which are the PDF of the density, the perimeter length $W_{1}$ and the Euler characteristic $\chi$. The latter two are defined as follows
@@
$\times\frac{\partial}{\partial\lambda_{\rho}}\left[\frac{\partial\varphi(\lambda_{\rho},\lambda_{g})}{\partial\lambda_{g}}\exp\left(\varphi(\lambda_{\rho},\lambda_{g})\right)\right]\,.$ (30)
@@ 4 Results
The Lévy flights are generated starting from an initial point in the center of a box of area $L\times L$. The flight is moved a distance $\ell$ from the previous point, in a direction drawn from a uniform angle $0\leq\theta<2\pi$. The $i^{\rm th}$ point in the flight is given by $\bf{x}_{i}={\bf x}_{i-1}+(\ell_{i}\cos\theta_{i},\ell_{i}\sin\theta_{i})$. Periodic boundary conditions are enforced, and ${\bf x}_{i}$ is adjusted modulo $L$ if the flight length is greater than $L$. Each point in the flight is binned into a regular lattice using a nearest-neighbour scheme, and we use resolution $\Delta=0.25$ and number of pixels per side in the box $N_{\rm pix}=4000$, so that the box size is $L=N_{\rm pix}\times\Delta$. We fix the number of Lévy flight steps as $N_{\rm pts}=40\times 2000^{2}$, $\alpha=3/2$ and vary $\ell_{0}$ to obtain fields with different $\xi_{0}$ values. Once the Lévy flight is binned into a regular lattice, we fast Fourier transform the field, apply a Gaussian smoothing kernel $W(kR)=e^{-k^{2}R^{2}/2}$, and inverse transform. We fix $R=1$ to match the window function used in the previous section. The particular case $\alpha=3/2$ is studied throughout this work, as it is a numerically tractable example of a Levy flight. Increasing $\alpha$ increases the large scale correlations of the points and generates larger finite box effects, whereas smaller $\alpha$ decreases the typical flight length and requires an increasingly large number of steps to populate the simulations.
@@
Before presenting the results, we first comment on a numerical systematic which hinders our ability to exactly reproduce the theoretical expectation values constructed in the previous section. In Figure 2 (top panel) we present the two-point correlation function of a set of $N_{\rm real}=100$ numerically generated Lévy flights with $\alpha=3/2$ and $\ell_{0}=0.02$ (red points/error bars). The expected large separation behaviour of the flight is $\zeta(r)\sim r^{\alpha-D}$ with $D=2$; this expected behaviour is presented as a blue dashed line. However, we find that the simulated Lévy flights present slightly steeper power law behaviour (solid blue line), corresponding to an effective $\alpha=1.43$ [^4] For larger separations, the correlation function drops faster than a power law and becomes increasingly noisy. In Figure 2 we present only the scales that dominate the statistical properties of the flights, which are well approximated by a power law.. This is a finite box effect; each step in the flight must be smaller than $\sim L$ due to the imposition of periodic boundary conditions. This leads to an excess probability of finding flight points at smaller separations, and leads to a systematic difference between theoretical expectation values and our numerical results [^5] Or alternatively one could model finite volume effects by replacing in equation (2) the integral over $\mathbf{k}$ by a sum over discrete values.. The problem will be exacerbated for larger values of $\alpha$, as the correlation function should become flat as $\alpha\to D$, which is a scenario that cannot be realised for a finite box. In what follows, for all summary statistics we present the expected theoretical results based on $\alpha=3/2$ and the effective result assuming $\alpha=1.43$. For mean field predictions, there is no $\alpha$ dependence.
@@ 5 Conclusion
Our results highlight the versatility of Rayleigh-Lévy flights, offering insights not only for astrophysical contexts but also for broader interdisciplinary applications. Rayleigh-Lévy flights model is a unique example of a random process that can both easily be implemented and whose statistical properties are explicitly known (in terms of correlation functions). Furthermore those correlation functions share qualitatively the same properties as those expected for cosmic density fields in both its quasi linear regime and its strongly non linear regime. This is therefore is very precious toy model for which analytical methods or numerical tools can be tested, from departure from Gaussian statistics to the exploration of the fully nonlinear regime. For example, one can obtain insights into the validity range of the Edgeworth expansion, which is commonly used to model the perturbative non-Gaussianity of the matter density of the Universe as traced by galaxies. The Edgeworth expansion is not a convergent series, so understanding what variances and density ranges we can use it to model the cosmological density field is an open question. The paper illustrates here that fully nonlinear objects such as peaks, critical points, and their correlations, can also be investigated in such a model, challenging and extending the picture that can be drawn from Gaussian fields. Furthermore writing numerical codes to evaluate the number of peaks, droughts, critical points in a discrete set of strongly correlated points can also be challenging. Having a toy model for which exact results are known is therefore very useful.